2. At the cose of the year a company has mventory ot S 150,000 and cost of goods solkd tor


$ 975,000. If the company's turnover ratio SS, detemine the opening balance of the mventory

1
Expert's answer
2021-11-14T17:43:33-0500

Solution:

The company’s inventory turnover ratio = 2.5

 

The formula for inventory turnover ratio = "\frac{Cost \;of\; goods \;sold}{Average\; inventory}"


Average inventory = (Opening inventory + Closing inventory) "\div" 2

Let opening inventory be = x

"2.5 = \frac{975,000}{(x+150,000)\div 2}"


"2.5 = \frac{975,000}{0.5x+75,000}"

Multiply both sides by 0.5x + 75000:

1.25x + 187,500 = 975,000

1.25x = 975,000 – 187,500

1.25x = 787,500

X = 630,000


Opening inventory = 630,000